OpenStudy (anonymous):

Someone help please! Will give medal! (On a time limit :( ) Given the function g(x) = -3x + 4, compare and contrast g(-2) and g(4). Choose the statement that is true concerning these two values. A. The value of g(-2) is smaller than the value of g(4). B. The value of g(-2) is the same as the value of g(4). C. The values of g(-2) and g(4) cannot be compared. D. The value of g(-2) is larger than the value of g(4).

OpenStudy (nurali):

If g(x)=-3x+4, then g(-2) = -3(-2)+4=6+4=10 g(4) = -3(4)+4=-12+4=-8 Since 10 > -21 The value of g(-2) is larger than the value of g(4).

OpenStudy (anonymous):

Thank you so much for explaining that to me!